# Active particle motion in Poiseuille flow through rectangular channels

### v1

@@ III Equilibrium states and stability
We start by finding equilibrium states of the dynamical system that would correspond to an active particle with a fixed cross-sectional location and a fixed orientation. This is done by making the time derivatives zero in Eq. (9) and solving the resulting nonlinear algebraic equations. We get the following two equilibrium states:
$(x^{*},y^{*},\theta^{*},\phi^{*})=(0,0,0,0)\,\,\&\,\,(0,0,0,\pi).$
The former equilibrium state, $(x^{*},y^{*},\theta^{*},\phi^{*})=(0,0,0,0)$, corresponds to an active particle oriented upstream at the center of the channel, while the latter state, $(x^{*},y^{*},\theta^{*},\phi^{*})=(0,0,0,\pi)$, corresponds to an active particle oriented downstream at the center of the channel.
To understand the stability of these equilibrium states, we perform a linear stability analysis by perturbing the equilibrium states as follows: $(x,y,\theta,\phi)=(x^{*},y^{*},\theta^{*},\phi^{*})+\epsilon(x_{1},y_{1},\theta_{1},\phi_{1})$, where $\epsilon>0$ is a perturbation parameter. Substituting this in Eq. (9) and comparing $O(\epsilon)$ terms we get a matrix equation for the evolution of perturbation variables $(x_{1},y_{1},\theta_{1},\phi_{1})$.
@@ III.0.1 Upstream-oriented equilibrium state
For the upstream-oriented equilibrium state $(x^{*},y^{*},\theta^{*},\phi^{*})=(0,0,0,0)$, we obtain the following matrix equation that governs the evolution of perturbations:
$\begin{bmatrix}\dot{x}_{1}\\ \dot{y}_{1}\\ \dot{\theta}_{1}\\ \dot{\phi}_{1}\end{bmatrix}=\begin{bmatrix}0&0&0&-1\\ 0&0&1&0\\ 0&-\bar{U}&0&0\\ \bar{U}/AR^{2}&0&0&0\end{bmatrix}\begin{bmatrix}{x}_{1}\\ {y}_{1}\\ {\theta}_{1}\\ {\phi}_{1}\\ \end{bmatrix}.$
@@
The stability of the equilibrium state is determined by the nature of the eigenvalues of the right-hand-side matrix [25]. We get the following characteristic polynomial equation for eigenvalues $\lambda$:
$\lambda^{4}+\bar{U}\left(1+\frac{1}{AR^{2}}\right)\lambda^{2}+\frac{\bar{U}^{2}}{AR^{2}}.$
The roots of this quartic polynomial gives the eigenvalues:
$\lambda=\pm i\sqrt{\bar{U}},\>\pm i\frac{\sqrt{\bar{U}}}{AR}.$
@@
Thus, all the eigenvalues are purely imaginary for this upstream-oriented active particle state. This equilibrium state may correspond to a center or a stable/unstable spiral [25], however, the conservative nature of our dynamical system makes this equilibrium point a center and we will revisit this numerically in Sec. V.3. Furthermore, due to the aspect ratio of the channel differing from $AR=1$ for non-square cross-sections, the oscillation frequency coming from the imaginary part of the eigenvalues is different in directions corresponding to the eigenvectors.
@@
Figure 2: Classification of active particle motion. Different trajectories of the active particle in a straight channel with square cross-section are shown in the $(z,y)$ plane and $(x,y)$ plane based on different initial positions for: (a) $(x(0),y(0))=(0.15,0.15)$, (b) $(x(0),y(0))=(0.05,0.30)$, (c) $(x(0),y(0))=(0.30,0.05)$, (d) $(x(0),y(0))=(0.58,0.70)$, (e) $(x(0),y(0))=(0.75,0.75)$ and (f) $(x(0),y(0))=(0.55,0.40)$. The different colors represent the classification of trajectory based on the region occupied in the cross-section. (a) Central swinging motion (green) - trajectories that stay near the center of the channel inside the gray square $-0.25<x,y<0.25$, (b) vertical swinging motion (cyan) - trajectories that are confined in the $x$ direction i.e. inside the vertical band $-0.25<x<0.25$, (c) horizontal swinging motion (blue) - trajectories that are confined in the $y$ direction i.e. inside the horizontal band $-0.25<y<0.25$, (d) off-centered trapping (yellow) - trajectories that are confined in a rectangular of area less than half the area of cross-section, and, the box crosses at most one centerline: the vertical centerline ($x=0$) or the horizontal centerline ($y=0$), (e) tumbling motion (purple) - trajectories that stay near the edges of the channel i.e. outside the region of central swinging motion, and (f) wandering motion (red) - trajectories that explore both the central as well as outer region of the cross-section. Other parameters were fixed to $\bar{U}=10$, $z(0)=0$, $\theta(0)=0$ and $\phi(0)=0$. See also supplemental videos S1-S6 for videos of trajectories corresponding to panels (a)-(f), respectively.
@@ III.0.2 Downstream-oriented equilibrium state
For the downstream-oriented equilibrium state $(x^{*},y^{*},\theta^{*},\phi^{*})=(0,0,0,\pi)$, we get the following matrix equation for evolution of perturbations:
$\begin{bmatrix}\dot{x}_{1}\\ \dot{y}_{1}\\ \dot{\theta}_{1}\\ \dot{\phi}_{1}\end{bmatrix}=\begin{bmatrix}0&0&0&1\\ 0&0&1&0\\ 0&\bar{U}&0&0\\ \bar{U}/AR^{2}&0&0&0\end{bmatrix}\begin{bmatrix}{x}_{1}\\ {y}_{1}\\ {\theta}_{1}\\ {\phi}_{1}\\ \end{bmatrix},$
with the following characteristic equation,
$\lambda^{4}-\bar{U}\left(1+\frac{1}{AR^{2}}\right)\lambda^{2}+\frac{\bar{U}^{2}}{AR^{2}}.$
@@
$\lambda=\pm\sqrt{\bar{U}},\>\pm\frac{\sqrt{\bar{U}}}{AR}.$
Since the eigenvalues are all real and both positive and negative, this equilibrium state is a saddle point having both stable and unstable manifolds. Hence this state is unstable. Again, due to the different magnitude of eigenvalues for non-square cross-sections, the rate of instability from the saddle point differs in the directions of the corresponding eigenvectors. We will numerically revisit the dynamics of an active particle starting near this equilibrium state in Sec. V.3.
@@ IV Dynamics in a cylindrical channel
Zöttl and Stark [16] studied the nonlinear dynamics of an active particle suspended in fluid flow through a cylindrical channel. For a circular cross-section, the flow field is axisymmetric and takes the form (in dimensionless form with length variables scaled with the duct radius)
$\bar{u}_{c}(x,y)=\bar{U}\left(1-(x^{2}+y^{2})\right).$ (10)
For this system, they identified two constants of motion as follows:
$H_{c}=\frac{1}{2}\bar{U}(x^{2}+y^{2})+1-\cos\phi\cos\theta,$ (11)
and
$L_{z}=x\sin\theta+y\cos\theta\sin\phi.$ (12)
@@ V Dynamics in a square channel
We now explore in detail the dynamics of an active particle suspended in fluid flow through a $3$D straight channel with a square cross-section i.e. $AR=1$. We explore the active particle dynamics as a function of the dimensionless parameter $\bar{U}$ as well as the initial conditions i.e. the initial position in the cross-section $(x(0),y(0))$, and the initial orientation angles $(\theta(0),\phi(0))$.
@@ V.1 Classification of active particle trajectories
A large diversity of trajectories are observed for the active particle in a square channel by varying the system parameter $\bar{U}$ as well as the initial conditions. Some typical trajectories and the corresponding orientations are shown in Fig. 2 for $\bar{U}=10$. We choose to classify the trajectories into the following six types based on the region they occupy in the square cross-section (gray region in Fig. 2): (i) Central swinging motion (green) - trajectories undergoing swinging motion about the channel centerline (similar to swinging motion in cylindrical channel [16]) and confined near the center of the channel. These trajectories are confined in the cross-sectional domain $-0.25<x,y<0.25$. (ii) Vertical swinging motion (cyan) - trajectories undergoing swinging motion in the vertical $y$ direction and confined in the $x$ direction. These trajectories are confined in a vertical band with $-0.25<x<0.25$. (iii) Horizontal swinging motion (blue) - trajectories undergoing swinging motion in the horizontal $x$ direction and confined in the $y$ direction. These trajectories occupy a horizontal band with $-0.25<y<0.25$. (iv) Off-centered trapping (yellow) - trajectories that are confined in a rectangular box of area less than half the area of the cross-section, and this box crosses at most one centerline but not both, i.e. the box crosses either the vertical centerline ($x=0$) or the horizontal centerline $(y=0)$ or none. (v) Tumbling motion (purple) - trajectories that stay outside the central region and wander near the channel walls. These trajectories occupy the region in the square cross-section which is outside the central region defined for central swinging motion. (vi) Wandering motion (red) - trajectories that visits both the central region defined for central swinging motion, as well as the outer region defined for tumbling motion.
@@ V.2 Comparison with cylindrical channel
For a square cross-section ($AR=1$) we have the approximate dimensionless fluid flow field from Eq. (8),
$\bar{u}_{s}(x,y)=\bar{U}(1-x^{2})(1-y^{2})=\bar{U}\left(1-(x^{2}+y^{2})+x^{2}y^{2}\right).$
By introducing the following flow field
$\bar{u}_{cs}(x,y)=\bar{U}\left(1-(x^{2}+y^{2})+\alpha x^{2}y^{2}\right),$
@@
we can continuously transform from a circular cross-section flow field in Eq. (10) $(\alpha=0)$ to the square cross-section flow field $\bar{u}_{s}(x,y)$ $(\alpha=1)$. The corresponding nonlinear ODEs for flow field $\bar{u}_{cs}(x,y)$ are:
$\dot{x}$ $=-\cos\theta\sin\phi$ (13)
$\dot{y}$ $=\sin\theta$
$\dot{\theta}$ $=-\bar{U}y\cos\phi+\alpha\bar{U}x^{2}y\cos\phi$
$\dot{\phi}$ $=\bar{U}\left(-y\tan\theta\sin\phi+x\right)+\alpha\bar{U}xy\left(x\tan\theta\sin\phi-y\right).$
@@
It can be seen from Eq. (13) that the equilibrium states and their linear stability found for a square cross-section ($AR=1$) in Sec. III will be the same as that for a circular cross-section Zöttl and Stark [16]. This is to be expected since the additional term present in the $\dot{\theta}$ and $\dot{\phi}$ equations in (13) contributes at an order $O(\epsilon^{3})$ near the equilibrium points and hence their effects are not felt at leading order i.e. $O(\epsilon)$.
For our square channel, we only have the constant of motion in Eq. (7) as
$H_{s}=\frac{1}{2}\bar{U}(x^{2}+y^{2}-x^{2}y^{2})+1-\cos\phi\cos\theta.$
@@
We chose $H_{s}=1+\frac{1}{2}\bar{U}+H_{g}\geq 0$, to better compare with $H_{c}$ in Eq. (11) for a circular cross-section. Furthermore, the time derivative of the second constant of motion $L_{z}$ for the cylindrical channel in Eq. (12), takes the following form for a square cross-section ($\alpha=1$):
$\frac{dL_{z}}{dt}=\bar{U}xy\cos\theta\cos\phi\left(x^{2}-y^{2}\right).$ (14)
@@
We see that this quantity will not be varying significantly when either $x$ or $y$ or both are small i.e. the motion is confined either near the center of the channel or along the $x$ or the $y$ axis. Hence, in this regime, we expect the motion of the system to be periodic/quasiperiodic since the system is close to the cylindrical channel system which is both Hamiltonian and integrable. Furthermore, from Eq. (14), the right-hand side will be near zero if the orientation is confined near $\theta=\pm\pi/2$ or $\phi=\pm\pi/2$ or if the motion is along diagonals $y=\pm x$. We expect our square cross-section system to deviate from a Hamiltonian integrable system when the motion takes place away from the channel center and chaos may arise. Hence, for swinging motion in the square cross-section which is confined near the center of the channel or along $x$ or $y$ axis, we typically expect regular (periodic/quasiperiodic) motion. Conversely, for tumbling motion that takes place away from the channel center, chaotic motion is possible. An example showing a comparison of active particle dynamics between circular and square cross-section for a typical tumbling motion with the same initial conditions $(x(0),y(0),\theta(0),\phi(0))=(0.55,0.58,\pi/24,\pi/24)$ is shown in supplemental video S7.
@@
Figure 3: Active particle motion for small perturbations near the unstable saddle equilibrium. Different trajectories of the active particles in the $(z,y)$ plane and $(x,y)$ plane for small perturbations: (a) in the direction of the unstable manifold $(x(0),y(0),\theta(0),\phi(0))=(0,0,0,\pi)+(0.02,0.02,0.02\sqrt{\bar{U}},0.02\sqrt{\bar{U}})$, (b) in the direction of the stable manifold $(x(0),y(0),\theta(0),\phi(0))=(0,0,0,\pi)+(0.02,0.02,-0.02\sqrt{\bar{U}},-0.02\sqrt{\bar{U}})$ and (c) small general perturbations $(x(0),y(0),\theta(0),\phi(0))=(0,0,0,\pi)+(0.01,0.02,0.015,-0.005)$. The parameter $\bar{U}=10$ was fixed and $z(0)=0$. See also supplemental videos S8-S10 for videos of panels (a)-(c), respectively.

### v2

@@ III Equilibrium states and stability
We start by finding equilibrium states of the dynamical system that would correspond to an active particle with a fixed cross-sectional location and a fixed orientation. This is done by making the time derivatives zero in Eq. (8) and solving the resulting nonlinear algebraic equations. We find the following two equilibrium states:
$(x^{*},y^{*},\theta^{*},\phi^{*})=(0,0,0,0)\,\,\text{and}\,\,(0,0,0,\pi).$
The first equilibrium (with $\phi^{*}=0$) corresponds to an active particle oriented upstream at the center of the channel, while the second (with $\phi^{*}=\pi$) corresponds to an active particle oriented downstream, also at the center of the channel.
To understand the stability of these equilibrium states, we perform a linear stability analysis, perturbing the equilibrium states thus: $(x,y,\theta,\phi)=(x^{*},y^{*},\theta^{*},\phi^{*})+\epsilon(x_{1},y_{1},\theta_{1},\phi_{1})$, where $0<\epsilon\ll 1$ is a perturbation parameter. Substituting this in Eq. (8) and comparing $O(\epsilon)$ terms we get a matrix equation for the evolution of the perturbation variables $(x_{1},y_{1},\theta_{1},\phi_{1})$.
@@ III.0.1 Upstream-oriented equilibrium state
For the upstream-oriented equilibrium state $(x^{*},y^{*},\theta^{*},\phi^{*})=(0,0,0,0)$, we obtain the following linear equation that governs the evolution of perturbations:
$\begin{bmatrix}\dot{x}_{1}\\ \dot{y}_{1}\\ \dot{\theta}_{1}\\ \dot{\phi}_{1}\end{bmatrix}=\begin{bmatrix}0&0&0&-1\\ 0&0&1&0\\ 0&-{U}&0&0\\ {U}/AR^{2}&0&0&0\end{bmatrix}\begin{bmatrix}{x}_{1}\\ {y}_{1}\\ {\theta}_{1}\\ {\phi}_{1}\\ \end{bmatrix}.$
@@
The stability of the equilibrium state is determined by the nature of the eigenvalues of the right-hand-side matrix [25]. We obtain the following characteristic polynomial equation for eigenvalues $\lambda$:
$\lambda^{4}+{U}\left(1+\frac{1}{AR^{2}}\right)\lambda^{2}+\frac{{U}^{2}}{AR^{2}}=0.$
The roots of this quartic polynomial give the eigenvalues
$\lambda=\pm i\sqrt{{U}},\>\pm i\frac{\sqrt{{U}}}{AR}.$
@@
All the eigenvalues being purely imaginary, this upstream-oriented active particle equilibrium may correspond to a center or a stable/unstable spiral [25]. However, the conservative nature of our dynamical system makes this equilibrium a center. We will revisit this numerically in Sec. V.3. Furthermore, for non-square cross-sections having aspect ratio differing from $AR=1$, the oscillation frequency differs along the two eigenvector pairs corresponding to the two conjugate eigenvalue pairs.
@@
Figure 2: Classification of active particle motion. Different trajectories of the active particle in a straight channel with square cross-section are shown in the $(z,y)$ plane and $(x,y)$ plane based on different initial positions for: (a) $(x(0),y(0))=(0.15,0.15)$, (b) $(x(0),y(0))=(0.05,0.30)$, (c) $(x(0),y(0))=(0.30,0.05)$, (d) $(x(0),y(0))=(0.58,0.70)$, (e) $(x(0),y(0))=(0.75,0.75)$ and (f) $(x(0),y(0))=(0.55,0.40)$. At right, the shaded region of the spherical surface shows the time evolution of the particle’s orientation. The different colors represent the classification of trajectory based on the region occupied in the cross-section. (a) Central swinging motion (green); trajectories that stay near the center of the channel inside the gray square $-0.25<x,y<0.25$. (b) Vertical swinging motion (cyan); trajectories that are confined in the $x$ direction i.e. inside the vertical band $-0.25<x<0.25$. (c) Horizontal swinging motion (blue); trajectories that are confined in the $y$ direction i.e. inside the horizontal band $-0.25<y<0.25$. (d) Off-centered trapping (yellow); trajectories that are confined in a rectangular region of the cross section having an area less than half the area of the cross-section and which may cross either centreline ($x=0$, $y=0$), or neither, but not both. (e) Tumbling motion (purple); trajectories that stay near the walls of the channel, i.e. outside the region of central swinging motion. (f) Wandering motion (red); trajectories that explore both the central as well as outer regions of the cross-section. Other parameters were fixed to ${U}=10$, $z(0)=0$, $\theta(0)=0$ and $\phi(0)=0$. See also supplemental videos S1-S6 for videos of particle trajectories and orientations corresponding to panels (a)-(f), respectively.
@@ III.0.2 Downstream-oriented equilibrium state
For the downstream-oriented equilibrium state $(x^{*},y^{*},\theta^{*},\phi^{*})=(0,0,0,\pi)$, we obtain the following linear equation describing the evolution of perturbations:
$\begin{bmatrix}\dot{x}_{1}\\ \dot{y}_{1}\\ \dot{\theta}_{1}\\ \dot{\phi}_{1}\end{bmatrix}=\begin{bmatrix}0&0&0&1\\ 0&0&1&0\\ 0&{U}&0&0\\ {U}/AR^{2}&0&0&0\end{bmatrix}\begin{bmatrix}{x}_{1}\\ {y}_{1}\\ {\theta}_{1}\\ {\phi}_{1}\\ \end{bmatrix},$
with the characteristic equation
$\lambda^{4}-{U}\left(1+\frac{1}{AR^{2}}\right)\lambda^{2}+\frac{{U}^{2}}{AR^{2}}=0.$
@@
$\lambda=\pm\sqrt{{U}},\>\pm\frac{\sqrt{{U}}}{AR}.$
Since the eigenvalues are all real, with two positive and two negative, this equilibrium state is an unstable saddle point having both stable and unstable manifolds which are two dimensional. Again, due to the different magnitudes of eigenvalues for non-square cross-sections, the rate of instability from the saddle point differs in the directions of the corresponding eigenvectors. We will numerically revisit the dynamics of an active particle starting near this equilibrium state in Sec. V.3.
@@ IV Dynamics in a cylindrical channel
Zöttl and Stark [16] studied the nonlinear dynamics of an active particle suspended in fluid flow through a cylindrical channel. For a circular cross-section, the flow field is axisymmetric and takes the dimensionless form
$u_{c}(x,y)={U}\left(1-(x^{2}+y^{2})\right),$ (9)
(with length variables scaled with the duct radius). For this special case, they identified two constants of motion:
$H_{c}$ $=\frac{1}{2}{U}(x^{2}+y^{2})+1-\cos\phi\cos\theta,$ (10a)
$L_{z}$ $=x\sin\theta+y\cos\theta\sin\phi.$ (10b)
@@ V Dynamics in a square channel
We now explore in detail the dynamics of an active particle suspended in fluid flow through a $3$D straight channel with a square cross-section i.e. $AR=1$. We explore the active particle dynamics as a function of the dimensionless parameter ${U}$ as well as the initial conditions i.e. the initial position in the cross-section $(x(0),y(0))$, and the initial orientation angles $(\theta(0),\phi(0))$.
@@ V.1 Classification of active particle trajectories
A large diversity of trajectories are observed for the active particle in a square channel by varying the system parameter ${U}$ as well as the initial conditions. Some typical trajectories and the corresponding orientations are shown in Fig. 2 for ${U}=10$. We choose to classify the trajectories into the following six types based on the region they occupy (shown in gray in Fig. 2) at long times within the square cross-section. (i) Central swinging motion (green) with trajectories undergoing swinging motion about the channel centerline (similar to swinging motion in cylindrical channel [16]) and confined near the center of the channel to the cross-sectional domain $-0.25<x,y<0.25$. (ii) Vertical swinging motion (cyan) with trajectories undergoing swinging motion in the vertical $y$ direction and confined in the $x$ direction to the vertical band $-0.25<x<0.25$. (iii) Horizontal swinging motion (blue) with trajectories undergoing swinging motion in the horizontal $x$ direction and confined in the $y$ direction to the horizontal band $-0.25<y<0.25$. (iv) Off-centered trapping (yellow) with trajectories confined within a rectangular region of the cross-section having an area less than half the area of the cross-section; this region may cross at most one centerline of the cross-section ($x=0$ or $y=0$ or neither) but not both. (v) Tumbling motion (purple) with trajectories that stay outside the central region of the cross-section as defined for central swinging, and wander near the channel walls. (vi) Wandering motion (red) with trajectories not in classes (i)–(v) and which, therefore, visit both the central region defined for central swinging motion, as well as the outer region defined for tumbling motion. Numerically, the classification is implemented by only analyzing the latter half of the trajectory to remove any transient dynamical behaviors at short times.
@@ V.2 Comparison with cylindrical channel
For a square cross-section ($AR=1$) we have the approximate dimensionless fluid velocity from Eq. (7),
${u}_{s}(x,y)={U}(1-x^{2})(1-y^{2})={U}\left(1-(x^{2}+y^{2})+x^{2}y^{2}\right).$
By introducing the following velocity field
${u}_{cs}(x,y)={U}\left(1-(x^{2}+y^{2})+\alpha x^{2}y^{2}\right),\ \alpha\in[0,1],$
@@
we can continuously transform from the Poiseuille flow of a circular cross-section as in Eq. (9) $(\alpha=0)$ to the approximate Poiseuille flow of a square cross-section ${u}_{s}(x,y)$ $(\alpha=1)$. The corresponding nonlinear ODEs for an active particle within the flow field ${u}_{cs}(x,y)$ are:
$\dot{x}$ $=-\cos\theta\sin\phi$ (11a)
$\dot{y}$ $=\sin\theta$ (11b)
$\dot{\theta}$ $=-{U}y\cos\phi+\alpha{U}x^{2}y\cos\phi$ (11c)
$\dot{\phi}$ $={U}\left(-y\tan\theta\sin\phi+x\right)$ (11d)
$\quad+\alpha{U}xy\left(x\tan\theta\sin\phi-y\right).$
@@
It can be seen from Eq. (11) that the equilibrium states and their linear stability, are the same for a square cross-section (refer Sec. III for $AR=1$) and a circular cross-section [16]. Specifically, since the $\alpha$ terms present in the $\dot{\theta}$ and $\dot{\phi}$ components of (11) contribute at an order $O(\epsilon^{3})$ near the equilibrium points their effects are not felt at order $O(\epsilon)$.
For our square channel, we only have the constant of motion $H_{g}$ in Eq. (6) (i.e. having implicitly utilised Eq. (3)). For ease of comparison with the quantity used by Zöttl and Stark [16] for a circular cross-section, $H_{c}$ of Eq. (10a), we define $H_{s}=1+\frac{1}{2}U+H_{g}\geq 0$, and express the constant of motion here as
$H_{s}=\frac{1}{2}{U}(x^{2}+y^{2}-x^{2}y^{2})+1-\cos\phi\cos\theta.$
@@
Furthermore, the time derivative of the second constant of motion $L_{z}$ for the cylindrical channel in Eq. (10b), has the following form for a square cross-section ($\alpha=1$):
$\frac{dL_{z}}{dt}={U}xy\cos\theta\cos\phi\left(x^{2}-y^{2}\right).$ (12)
@@
We see that this quantity will not vary significantly when either $x$ or $y$ (or both) are small, or when $y\approx\pm x$. Thus, for motion in a square cross-section confined near the channel center, along $x$ or $y$ axis or along diagonals, we expect the dynamics of the system to be regular (periodic/quasi-periodic) since the system is close to the cylindrical channel system which is both Hamiltonian and integrable. The central, vertical and horizontal swinging motions shown in Fig. 2 are examples of this type of motion. Conversely, for general motion that is not restricted to these above regions such that the particle explores regions away from the center of the channel cross-section, the system deviates from a Hamiltonian integrable system and chaotic motion may arise. Tumbling and wandering motions (see Fig. 2) are examples of this. A comparison of active particle dynamics between circular and square cross-sections for a typical tumbling motion with the same initial conditions $(x(0),y(0),\theta(0),\phi(0))=(0.55,0.58,\pi/24,\pi/24)$ is shown in supplemental video S7.
@@
Figure 3: Active particle motion for small perturbations near the unstable saddle equilibrium. Different trajectories are shown in the $(z,y)$ plane and $(x,y)$ plane for small perturbations (a) in the direction of the unstable manifold $(x(0),y(0),\theta(0),\phi(0))=(0,0,0,\pi)+(0.02,0.02,0.02\sqrt{{U}},0.02\sqrt{{U}})$, (b) in the direction of the stable manifold $(x(0),y(0),\theta(0),\phi(0))=(0,0,0,\pi)+(0.02,0.02,-0.02\sqrt{{U}},-0.02\sqrt{{U}})$ and (c) a small general perturbation $(x(0),y(0),\theta(0),\phi(0))=(0,0,0,\pi)+(0.01,0.02,0.015,-0.005)$. The parameter ${U}=10$ was fixed and $z(0)=0$. See also supplemental videos S8-S10 for videos of panels (a)-(c), respectively.